# Spectral Convolutional Conditional Neural Processes

### v5

@@ B.4.1 Model Architectures
Because the data lie on a regular grid, we use the on-the-grid implementation of the ConvCNP [gordon2019convolutional], eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote an incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the output of the convolutional deep set module has two components: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{y}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [gordon2019convolutional].
The density channel is obtained by convolving this modified filter with the mask $\mathrm{M}_{c}$. The kernel-smoothed component is computed by first multiplying $\mathrm{I}$ elementwise with $\mathrm{M}_{c}$ (setting non-context pixels to zero) and then applying the same modified convolution. Following gordon2019convolutional, we omit normalization by the density channel, as it did not provide empirical benefits and occasionally introduced instability; however, we retain the positivity constraint on the kernel.
The resulting representation is processed pointwise by an MLP with two hidden layers of width 128. These features are then fed into a ResNet-style CNN [he2016deep] consisting of six residual convolutional blocks (kernel size 11, 128 channels), following the implementation of bruinsma2023autoregressive. The embeddings corresponding to query pixels are finally passed to a decoder MLP with two hidden layers of width 128.
@@ B.4.2 Data and Experimental Setup
For our image-completion experiments, we use the Describable Textures Dataset (DTD; cimpoi2014describing). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then downsample to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid along each axis are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
Models are trained for 500 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. At each epoch, the model processes batches of 16 tasks. For each batch, we draw the number of context pixels as $|\mathcal{I}_{c}|\sim\mathcal{U}[5,1024)$, and treat all remaining pixels as queries, i.e. $|\mathcal{I}_{q}|=64\times 64-|\mathcal{I}_{c}|$. A single sampled value of $|\mathcal{I}_{c}|$ is shared across all tasks within the batch.

### v6

@@ B.4.1 Model Architectures
Because the data lie on a regular grid, we use the on-the-grid implementation of the ConvCNP [Gordon et al., 2019], eliminating the discretization step and improving computational efficiency. Let $\mathrm{I}$ denote an incomplete image, with unobserved pixels filled with dummy values, and let $\mathrm{M}_{c}$ be the binary mask indicating observed (context) pixels. For multi-channel images, the mask is broadcast along the channel dimension.
As in Section B.1.1, the output of the convolutional deep set module has two components: (i) a density channel capturing the spatial distribution of context pixels, and (ii) a kernel-smoothed representation. The kernel is implemented via a 2D convolutional layer (kernel size 11, $d_{y}$ input channels, 128 output channels, no bias) with a nonnegativity constraint enforced by taking absolute values of the learned weights during the forward pass [Gordon et al., 2019].
The density channel is obtained by convolving this modified filter with the mask $\mathrm{M}_{c}$. The kernel-smoothed component is computed by first multiplying $\mathrm{I}$ elementwise with $\mathrm{M}_{c}$ (setting non-context pixels to zero) and then applying the same modified convolution. Following Gordon et al. [2019], we omit normalization by the density channel, as it did not provide empirical benefits and occasionally introduced instability; however, we retain the positivity constraint on the kernel.
The resulting representation is processed pointwise by an MLP with two hidden layers of width 128. These features are then fed into a ResNet-style CNN [He et al., 2016] consisting of six residual convolutional blocks (kernel size 11, 128 channels), following the implementation of Bruinsma et al. [2023]. The embeddings corresponding to query pixels are finally passed to a decoder MLP with two hidden layers of width 128.
@@ B.4.2 Data and Experimental Setup
For our image-completion experiments, we use the Describable Textures Dataset (DTD; Cimpoi et al. [2014]). We adopt the standard DTD split, which contains 1,880 images each for training, validation, and testing. All images are RGB, with heights ranging from 231 to 778 pixels and widths from 271 to 900 pixels. During training, validation, and testing, we sample from each image a random $192\times 192$ crop, which we then downsample to $64\times 64$. Pixel coordinates of the resulting $64\times 64$ grid along each axis are linearly mapped to [1,1][-1,1], and pixel intensities are normalized to $[0,1]$ independently across channels.
Models are trained for 500 epochs using AdamW with a learning rate of $10^{-4}$. We apply gradient clipping with a maximum norm of $0.5$. At each epoch, the model processes batches of 16 tasks. For each batch, we draw the number of context pixels as $n_{c}\sim\mathcal{U}[5,1024)$, and treat all remaining pixels as queries, i.e. $n_{q}=64\times 64-n_{c}$. A single sampled value of $n_{c}$ is shared across all tasks within the batch.